[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jul 16, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Jul 16 09:23:06 UTC 2017


HSU Shift Log for Sun Jul 16 2017 05:22:57 GMT-0400 (EDT)

Shift: A

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: Lots of GOES-16 IR data dropouts on shift; however,  
no problems were noted with the other channels.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170716/4af9ca58/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list